Calculate the electric field strength of a proton experiencing an electric force of 3.75x10-14N. 


1
Expert's answer
2021-02-18T18:54:03-0500
"E=\dfrac{F}{q}=\dfrac{3.75\cdot10^{-14}\ N}{1.6\cdot10^{-19}\ C}=2.34\cdot10^5\ \dfrac{N}{m}."